The total of two number is 23. double the large number exceeds three times the smaller by sixteen. what are the integers.
Answer by checkley71(8403)

X+Y=23 OR X=23-Y
2Y=3X+16
NOW SUBSTITUTE (23-Y) FOR X & SOLVE FOR Y
2Y=3(23-Y)+16
2Y=69-3Y+16
2Y+3Y=85
5Y=85
Y=85/5
Y=17 ANSWER.
X+17=23
X=23-17
X=6 ANSWER.
PROOF
2*17=3*6+16
34=18=16
34=34